show work Matthews Fish Fry has a monthly target operating income of $7,900. Variable expenses are 70% of sales and monthly fixed expeSlamburgers sells four cheeseburgers for every two regular hamburgers. A cheeseburger sells for $2.75 with a variable cost o
Answer #1

Question 1

Correct answer-------$26,333

Working

A Contribution margin ratio 30%
B Fixed cost $          1,290
C=B/A Break even sales 4300

.

A Contribution margin ratio 30%
B Fixed cost+ Desired profit $          9,190
C=B/A Actual sales $        30,633

.

A Actual sales $        30,633
B Break even sales $          4,300
A-B Margin of safety $        26,333

.

Question 2

Correct answer-------$0.83

Working

Working Cheeseburger Hamburger Total
A Price $                        2.75 $                       2.00
B Variable Cost per unit $                        2.25 $                       0.50
C = A - B Contribution Margin $                        0.50 $                       1.50
D Product Mix 66.67% 33.33% 100.00%
E = C x D Weighted Average Contribution Margin $                        0.33 $                       0.50 $                   0.83